Are the grades one receives in a course related to the amount of time spent studying the subject? To analyze this possibility, a student took a random sample of 60 students who had enrolled in an accounting class last semester. She asked each student to report his or her grade in the course and the total number of hours spent studying accounting. The data are stored in this Excel file (SEE DATA BELOW)

Question 4. How many points would a student's course grade increase if the student studied an additional 5 hours? (Use 2 decimal places in your answer)

24 0 2 2 3 2 1 3 2 3 4 8 2 9 6 3 9 372505662427 3121272 23 3 2 3 3 4 3 8 3 3 3 3 1 2 3 123456789 10 1 2 3 4 5 6 7 8 9 20 21 22 23 24 25 26 27 28 29 3 3 3 3 34 3 3 3 3 39 40 41 42 43 44 45 46 4 4 49 50 553 54 5 56 7 8 9 60 61 62 63 64 65 66 67Question 5. If a students wants to increase her grade by 3 points, how many additional hours should she study? (Use 2 decimal places in your answer)
Answer #1

Answer:

Following is the output of regression analysis:

SUMMARY OUTPUT Regression Statistics MultipleR R Square Adjusted R Square Standard Error Observations 0.747524 0.558791 0.551

The equation of regression line:

y = 1.114*x +34.806

Question 4: How many points would a student's course grade increase if the student studied an additional 2 hours? (Use 2 decimal places in your answer)

Slope of line = 1.114

For each unit increase in study hours, student's course garde is increased by 1.114. So if the student studied an additional 2 hours, a student's course grade increased by 2 * 1.114 = 2.228

Answer: 2.23

Question 5:If a students wants to increase her grade by 3 points, how many additional hours should she study? (Use 2 decimal places in your answer)

Slope of line = 1.114

For each unit increase in study hours, student's course garde is increased by 1.114. Let x shows the required hours. We need x such that

x * 1.114 = 3

x = 2.69 hours

Thank you!